OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18

OpenStudy (anonymous):

x = 20-6y/13 , x= 18-4y/7 comparing 20-6y/13 = 18-4y/7 140- 42y=234-28y -28y+42y=140-234 14y= -94 y =-94/14 y=-6.71